OpenStudy (anonymous):

heeelp A local snail population grows according to the function g(x) = 200(1.03)2^x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent.

OpenStudy (sparklestaraa):

hhh :P lesson what? Module what? :P

OpenStudy (sparklestaraa):

Same problem b4 here :P @KaylaIsBae

OpenStudy (sparklestaraa):

Omg i've had such problems. Im at the end almost of my Part A but its confusing ur in LS right?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@mathslover @ganeshie8 please help

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

Delete all unnecessary comments..

OpenStudy (anonymous):

Is this right? \[g(x) = 200(1.03) 2^x\]

OpenStudy (anonymous):

I did

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry so much lag diving me mad

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@robtobey help?

OpenStudy (bohotness):

what is giving you trouble?

OpenStudy (anonymous):

i don't know how to do it

OpenStudy (anonymous):

@perl this is my question

OpenStudy (perl):

I think the question should be g(x) = 200(1.03)^(2x) now the question is saying , turn this into a function with only x as the exponent

OpenStudy (anonymous):

yeah and all i see is question marks ^

OpenStudy (perl):

g(x)=200(1.03)^(2x) =200(1.03^2)^x = 200(1.0609)^x

OpenStudy (anonymous):

yeah i see

OpenStudy (anonymous):

oooh, so the final answer is just 200(1.0609)^x or g(x) = 200(1.0609)^x

OpenStudy (perl):

correct :)

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

then it says Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. key features is like the principal and the rate of change, right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@perl ?